# ContextFlow: Training-Free Video Object Editing via Adaptive Context Enrichment

### v1

@@ 1 Introduction
Currently, there are two primary technical paths in the research community: training-based and training-free methods. (1) Training-based methods aim to build powerful and feed-forward models. Recent examples include video propagation-based models like I2V Edit [38], GenProp [22] and ReVideo [36], as well as other architectures such as VideoAnyDoor [54], GetIn [25], VACE [4] and UNIC [65], which achieve impressive results on respective benchmarks. However, these training-based methods are limited by the prohibitive computational costs and the demand for expensive large-scale datasets. (2) Compared with the training-based methods, the training-free method offers a more flexible and cost-effective alternative. Early works, like AnyV2V [18], leverage the vast knowledge embedded in pre-trained foundation models, obviating the need for any task-specific fine-tuning. This paradigm typically relies on a foundational workflow: first, inverting the source video to a noise latent using DDIM Inversion [49], and then guiding the new generation via Plug-and-Play (PnP) feature injection [10]. In this process, key internal features from a reconstruction of the source video are injected into the generation process of the edited video to enforce structural consistency.
@@
To address these critical limitations, we propose ContextFlow, a novel training-free framework that significantly advances the editing process for DiT-based models. Instead of relying on lossy inversion and crude feature replacement, ContextFlow establishes a high-fidelity, highly reversible foundation for editing based on RF-Solver [58]. At its core is a novel Adaptive Context Enrichment mechanism, which empowers the model to dynamically fuse information from the original video and the desired edit on a per-token basis. This dynamic approach addresses the conflict between content preservation and synthesis. To apply this guidance with efficiency and precision, we propose a systematic, data-driven Vital Layer Analysis to identify the most crucial intervention points within the DiT architecture.
@@ 2.1 Diffusion-based Video Editing
The landscape of video editing has been reshaped by diffusion models. Current research in this area is largely bifurcated into tuning-based and training-free methods. Tuning-based methods adapt pre-trained models for specific videos or subjects, exemplified by per-video optimization [62] and subject-driven personalization [69, 68, 8, 66, 64, 27, 46, 21, 35, 63, 61, 12, 13]. Other approaches focus on decoupling content and motion control during training to offer more flexible and precise editing capabilities [36, 33, 38, 29, 28, 27, 31, 3, 67, 39, 26]. While powerful, these methods often incur significant computational costs. Training-free methods aim for zero-shot video editing without model training. Early efforts focused on preserving video structure by manipulating cross-frame attention maps [57, 9, 7, 34, 32, 42]. A significant recent trend, however, has shifted towards propagating a modified first frame by leveraging the strong generative priors of pretrained Image-to-Video (I2V) models. Seminal works in this direction, such as AnyV2V [18], have demonstrated impressive consistency and quality. This paradigm has been further generalized in frameworks like UniEdit [1] and GenProp [22]. In parallel, other works [15, 5, 16, 30, 56, 8, 24, 2] have focused on customizing the diffusion process itself to enhance motion control and appearance preservation.
@@ 2.2 Reference-Guided Object Editing
A critical limitation of text-only guidance is its inability to specify the unique visual identity of a particular real-world object. This has spurred the development of reference-guided editing, which uses an image to define the target’s appearance. A common principle is establishing feature-level correspondence to guide attention and ensure faithful appearance propagation. Techniques like Zero-to-Hero [51] establish robust feature-level connections between the reference image and video frames, guiding attention mechanisms to ensure faithful appearance transfer. This principle of using correspondence to guide propagation is shared by a variety of recent works that aim for consistent reference-based editing, including VideoSwap [11], AnyV2V [18], and I2VEdit [38]. A more formidable challenge is to seamlessly insert a new object from a reference image into an existing video. Training-free methods such as InVi [47], GetInVideo [70], and iDiT-HOI [48] address demands that include not only identity preservation and temporal coherence but also plausible interaction with the scene. These editing-focused methods differ from reference-guided generation frameworks like VideoAnydoor [54] and MotionCtrl [60], which synthesize entirely new videos from scratch based on a reference object and motion prompts, rather than modifying pre-existing footage. Our approach aims to unify object-related video editing tasks, constructing a training-free framework based on the latest DiT model to simultaneously support multiple functions such as object insertion, object swapping, and object deletion.
@@ 3.3 High-Fidelity Inversion via Rectified Flow
Our editing workflow begins with the source video $V_{src}$ and a target prompt $s_{tgt}$. Using an off-the-shelf image editor (e.g., AnyDoor [6]), we first modify the initial frame $V_{src}[0]$ to create the edited frame $I_{edit}$. The primary challenge is then to propagate the static edit in $I_{edit}$ throughout the video, guided by $s_{tgt}$, while maintaining fidelity to $V_{src}$.
As established in our preliminaries, the success of this editing hinges on the quality of the initial noise latent. Standard inversion methods like DDIM are lossy, making it difficult to disentangle editing artifacts from inversion errors. To eliminate this ambiguity, a higher-order numerical solver is essential. Therefore, we introduce RF-Solver [58], a training-free, second-order sampler that provides the highly reversible and high-fidelity mapping required for a robust generative foundation.
@@ 3.5 Targeted Fusion via Vital Layer Analysis
A high $GR_{l}$ score signifies that the layer is highly sensitive to the guidance and thus influential in the editing process. Applying this analysis across a range of editing tasks reveals that layer responsiveness is not uniform, but instead exhibits a highly structured and task-dependent pattern, as shown in Figure 5. There are three primary zones of high responsiveness across the model’s depth: an early block (layers  1-10), a middle block (layers  15-21), and a deep block (layers  26-32). Crucially, the dominant zone of activity varies systematically with the task. For object insertion, peak responsiveness is consistently located in the early-layer block. In contrast, object swapping elicits the strongest response in the deep-layer block. Object deletion presents a unique dual-peak pattern, showing high responsiveness in both the middle and deep blocks. These empirical evidence strongly suggests a structural pattern within the DiT’s layers. The observed patterns are consistent with the broader understanding of Transformer architectures, where early layers typically handle spatial and structural information, while deeper layers manage more abstract semantic concepts [53, 44]. For instance, the reliance of insertion on early layers aligns with a need to establish spatial layout, while deletion’s dependence on deep layers corresponds to a high-level semantic operation.
@@ 4.1 Experimental Setup
Our framework is built upon Wan2.1-I2V-14B-480P [55], a publicly available image-to-video Diffusion Transformer with 40 layers. We adhere to a training-free paradigm, requiring no optimization or fine-tuning of the pre-trained diffusion model. For inversion, we utilize RF-Solver [58] with 50 steps to map the source video into a noise latent, and the subsequent editing also uses 50 sampling steps. We set the timestep threshold $\tau$ to 0.5 (i.e., the mechanism is adopted for the first 50% of timesteps) and a guidance scale of 3.0.
@@
We evaluate our method on the Unic-Benchmark [65]. For first-frame editing, we use AnyDoor [6] for insertion, InsertAnything [50] for swapping, and MagicQuill [23] for deletion. We compare against baselines including training-free AnyV2V [18], our adapted AnyV2V-DiT, and training-based VACE [4] and I2VEdit [38]. For evaluation, we measure task-specific fidelity (CLIP-score, DINO-score), background preservation (PSNR, SSIM), and overall video quality using metrics from VBench [14] (e.g., consistency, smoothness, dynamic and aesthetic quality).
@@ 5 Conclusion
  • Additional Qualitative Comparisons. We present more visual comparisons against some state-of-the-art video editing techniques methods, specifically Unic [65] and Pika [40], to further highlight the superiority of ContextFlow.
@@ 6.1 Experimental Setup and Computational Cost
This section expands on the implementation details mentioned in the main paper. Our framework is implemented in Python using PyTorch. All experiments were conducted on a server with two NVIDIA A800 (80GB) GPUs. We build our training-free framework upon the publicly available Wan2.1-I2V-14B-480P [55] model. The dual-path generation process, which is central to our Adaptive Context Enrichment mechanism, requires approximately 120GB of VRAM in total ( 60GB per GPU). For a typical 81-frame video at 480p resolution, the dual-path generation process takes approximately 25 minutes. As our method propagates an edit from the first frame, we utilize off-the-shelf image editing tools to generate the initial edited frame $I_{edit}$. For the experiments presented, we use AnyDoor [6] for object insertion, InsertAnything [50] for object swapping, and MagicQuill [23] for object deletion.
@@ 6.2 Further Details on Guidance Responsiveness Analysis
Expanding on the Guidance Responsiveness analysis presented in Figure 5 of the main paper, we offer details on its robustness. The $GR$ curves were derived by averaging results from a diverse set of 10 videos for each task, including in-domain videos from the Unic-Benchmark [65] and out-of-domain videos from sources like the DAVIS dataset [41].
@@ 6.3 Evaluation Protocol Details
To measure how well the identity of the reference object is preserved, we compute frame-wise similarity. For CLIP-I, we use the pre-trained clip-vit-large-patch14 [43] model from OpenAI. For DINO-I, we use the dinov2-giant [37] model from Facebook. The process is as follows: (1) we extract the image feature of the reference object; (2) for each frame in the generated video, we extract its image feature; (3) we compute the cosine similarity between the reference feature and each frame’s feature; (4) the final score is the average similarity across all frames.
@@
These metrics are computed exclusively on the unedited regions of the video to measure background stability. The unedited area is defined by a mask. For object swapping and deletion tasks on the Unic-Benchmark, we use the ground-truth masks provided with the dataset. For the object insertion task, where no prior object exists, we generate a mask for the newly inserted object using Grounding-SAM [45], and the background is defined as the area outside this mask. The final PSNR and SSIM scores are averaged over all frames.
@@
For the reader’s convenience, we briefly summarize the calculation methods for the VBench metrics used in our paper, drawing from the descriptions in the official VBench suite [14].
  • Overall Consistency. This metric evaluates the overall video-text consistency, reflecting both semantic and style alignment. It is computed using ViCLIP [59] as an aiding metric, where the model assesses the video against text prompts that contain different semantics and styles.
  • Motion Smoothness. To evaluate temporal quality beyond simple appearance consistency, this metric assesses whether the motion in the generated video is smooth and follows the physical laws of the real world. It utilizes the motion priors from a video frame interpolation model [20] to quantify the smoothness of the generated motions.
  • Dynamic Degree. Complementing smoothness, this metric measures the degree of dynamics (i.e., the presence of large motions) to ensure that completely static videos are not unfairly favored in temporal quality scores. It uses the RAFT [52] model to estimate the degree of dynamics in synthesized videos.
  • Aesthetic Quality. This metric evaluates the frame-wise artistic and beauty value perceived by humans. It employs the LAION [19] aesthetic predictor to assess aspects such as the layout, the richness and harmony of colors, photo-realism, naturalness, and the overall artistic quality of the video frames.
@@ 7 Additional Qualitative Comparisons
To further situate ContextFlow’s performance, we provide qualitative comparisons against more strong, contemporary methods, for example Unic [65] and commercial training-based tools like Pika  [40] and Kling  [17]. As shown in Figure 7 - Figure 14 , ContextFlow consistently demonstrates superior performance in both identity preservation and motion consistency.
@@ 8 Visual Analysis for Ablation Studies
Figure 20 and Figure 21 validates our data-driven layer selection strategy. Injecting information into all layers will lead to an excessive amount of information from the reconstruction path being introduced into the editing path, making it difficult to maintain the edited objects and causing the generated video to unconsciously lean towards the original video. Injecting guidance into the four least-responsive layers yields poor results, similar to providing weak guidance. Using a U-Net-based heuristic (from AnyV2V [18]) in the DiT model, which in a sense amounts to randomly selecting certain layers within DiT for injection, still leads to suboptimal results. Due to the inability of the injected layers to solve substantive problems, it often leads to stiff interaction between new objects and the background, and even unconscious emergence of unreasonable phenomena (such as the chimpanzee in Figure 21 strangely growing a tail). This visually confirms that our Guidance Responsiveness metric correctly identifies the most impactful layers for intervention.
@@
Figure 13: Visualization of Qualitative Comparison. Target prompt: “A hand-held white plastic watering can is used to water a vibrant collection of pink and purple flowers, including geraniums and possibly petunias, in a white planter.” The target to be deleted in this video is the railing in the lower left corner. Our method has high fidelity for the unedited areas, and its effect is better than that of Kling  [17].
@@
Figure 20: Visualization of Ablation on Layer Selection. Target prompt: “A Pikachu is floating on the sea.” Injection on all layers and Injection on 4 least responsive layers will cause Pikachu to disappear quickly in subsequent frames or fail to display actions correctly according to the target prompt. Injection selection follow AnyV2V [18] will produce an extremely stiff and weird “floating” effect that is highly unnatural, along with an odd smile not required by the prompt. Please zoom in for a closer check.
Figure 21: Visualization of Ablation on Layer Selection. Target prompt: “A red poppy flower surrounded by purple flowers. A large gorilla is gently trying to touch the red poppy flower.” Injection on all layers and Injection on 4 least responsive layers will cause the inserted object to disappear or become translucent. Injection selection follow AnyV2V [18] will once again result in the strange form of the gorilla: an oddly appearing tail (within the red circle). Please zoom in for a closer check.

### v2

@@ 1 Introduction
Currently, there are two primary technical paths in the research community: training-based and training-free methods. (1) Training-based methods aim to build powerful and feed-forward models. Recent examples include video propagation-based models like I2V Edit [ouyang2024i2vedit], GenProp [liu2024generativevideopropagation] and ReVideo [mou2024revideo], as well as other architectures such as VideoAnyDoor [tu2025videoanydoor], GetIn [zhuang2025videoaddwantvideo], VACE [chen2024vace] and UNIC [ye2025unic], which achieve impressive results on respective benchmarks. However, these training-based methods are limited by the prohibitive computational costs and the demand for expensive large-scale datasets. (2) Compared with the training-based methods, the training-free method offers a more flexible and cost-effective alternative. Early works, like AnyV2V [ku2024anyv2v], leverage the vast knowledge embedded in pre-trained foundation models, obviating the need for any task-specific fine-tuning. This paradigm typically relies on a foundational workflow: first, inverting the source video to a noise latent using DDIM Inversion [song2020denoising], and then guiding the new generation via Plug-and-Play (PnP) feature injection [graikos2022diffusion]. In this process, key internal features from a reconstruction of the source video are injected into the generation process of the edited video to enforce structural consistency.
@@
To address these critical limitations, we propose ContextFlow, a novel training-free framework that significantly advances the editing process for DiT-based models. Instead of relying on lossy inversion and crude feature replacement, ContextFlow establishes a high-fidelity, highly reversible foundation for editing based on RF-Solver [wang2024taming]. At its core is a novel Adaptive Context Enrichment mechanism, which empowers the model to dynamically fuse information from the original video and the desired edit on a per-token basis. This dynamic approach addresses the conflict between content preservation and synthesis. To apply this guidance with efficiency and precision, we propose a systematic, data-driven Vital Layer Analysis to identify the most crucial intervention points within the DiT architecture.
@@ 2.1 Diffusion-based Video Editing
The landscape of video editing has been reshaped by diffusion models. Current research in this area is largely bifurcated into tuning-based and training-free methods. Tuning-based methods adapt pre-trained models for specific videos or subjects, exemplified by per-video optimization [wu2023tune] and subject-driven personalization [zhu2025multibooth, zhu2024instantswap, feng2025follow, zhang2025magiccolor, yan2025eedit, ma2022visual, ruiz2023dreambooth, liu2025avatarartist, molad2023dreamix, wu2024customcrafter, wei2024dreamvideo2, he2024id, he2025fulldit2]. Other approaches focus on decoupling content and motion control during training to offer more flexible and precise editing capabilities [mou2024revideo, ma2025followyourclick, ouyang2024i2vedit, ma2024followpose, ma2023magicstick, ma2022visual, ma2025controllable, che2024gamegen, zheng2024videogen, pang2024dreamdance, ma2025model]. While powerful, these methods often incur significant computational costs. Training-free methods aim for zero-shot video editing without model training. Early efforts focused on preserving video structure by manipulating cross-frame attention maps [wang2024cove, geyer2023tokenflow, feng2025dit4edit, ma2025followyourmotion, ma2025followcreation, qi2023fatezero]. A significant recent trend, however, has shifted towards propagating a modified first frame by leveraging the strong generative priors of pretrained Image-to-Video (I2V) models. Seminal works in this direction, such as AnyV2V [ku2024anyv2v], have demonstrated impressive consistency and quality. This paradigm has been further generalized in frameworks like UniEdit [bai2024uniedit] and GenProp [liu2024generativevideopropagation]. In parallel, other works [jeong2024vmc, chen2024follow, kara2024rave, ma2024followyouremoji, wang2025consistent, feng2025follow, long2025follow, burgert2025gowiththeflow] have focused on customizing the diffusion process itself to enhance motion control and appearance preservation.
@@ 2.2 Reference-Guided Object Editing
A critical limitation of text-only guidance is its inability to specify the unique visual identity of a particular real-world object. This has spurred the development of reference-guided editing, which uses an image to define the target’s appearance. A common principle is establishing feature-level correspondence to guide attention and ensure faithful appearance propagation. Techniques like Zero-to-Hero [su2025zerotohero] establish robust feature-level connections between the reference image and video frames, guiding attention mechanisms to ensure faithful appearance transfer. This principle of using correspondence to guide propagation is shared by a variety of recent works that aim for consistent reference-based editing, including VideoSwap [gu2024videoswap], AnyV2V [ku2024anyv2v], and I2VEdit [ouyang2024i2vedit]. A more formidable challenge is to seamlessly insert a new object from a reference image into an existing video. Training-free methods such as InVi [saini2024invi], GetInVideo [zhuang2025videoaddwantvideo], and iDiT-HOI [shen2025idithoi] address demands that include not only identity preservation and temporal coherence but also plausible interaction with the scene. These editing-focused methods differ from reference-guided generation frameworks like VideoAnydoor [tu2025videoanydoor] and MotionCtrl [wang2024motionctrl], which synthesize entirely new videos from scratch based on a reference object and motion prompts, rather than modifying pre-existing footage. Our approach aims to unify object-related video editing tasks, constructing a training-free framework based on the latest DiT model to simultaneously support multiple functions such as object insertion, object swapping, and object deletion.
@@ 3.3 High-Fidelity Inversion via Rectified Flow
Our editing workflow begins with the source video $V_{src}$ and a target prompt $s_{tgt}$. Using an off-the-shelf image editor (e.g., AnyDoor [chen2024anydoor]), we first modify the initial frame $V_{src}[0]$ to create the edited frame $I_{edit}$. The primary challenge is then to propagate the static edit in $I_{edit}$ throughout the video, guided by $s_{tgt}$, while maintaining fidelity to $V_{src}$.
As established in our preliminaries, the success of this editing hinges on the quality of the initial noise latent. Standard inversion methods like DDIM are lossy, making it difficult to disentangle editing artifacts from inversion errors. To eliminate this ambiguity, a higher-order numerical solver is essential. Therefore, we introduce RF-Solver [wang2024taming], a training-free, second-order sampler that provides the highly reversible and high-fidelity mapping required for a robust generative foundation.
@@ 3.5 Targeted Fusion via Vital Layer Analysis
A high $GR_{l}$ score signifies that the layer is highly sensitive to the guidance and thus influential in the editing process. Applying this analysis across a range of editing tasks reveals that layer responsiveness is not uniform, but instead exhibits a highly structured and task-dependent pattern, as shown in Figure 5. There are three primary zones of high responsiveness across the model’s depth: an early block (layers  1-10), a middle block (layers  15-21), and a deep block (layers  26-32). Crucially, the dominant zone of activity varies systematically with the task. For object insertion, peak responsiveness is consistently located in the early-layer block. In contrast, object swapping elicits the strongest response in the deep-layer block. Object deletion presents a unique dual-peak pattern, showing high responsiveness in both the middle and deep blocks. These empirical evidence strongly suggests a structural pattern within the DiT’s layers. The observed patterns are consistent with the broader understanding of Transformer architectures, where early layers typically handle spatial and structural information, while deeper layers manage more abstract semantic concepts [tenney2019bert, raghu2021vision]. For instance, the reliance of insertion on early layers aligns with a need to establish spatial layout, while deletion’s dependence on deep layers corresponds to a high-level semantic operation.
@@ 4.1 Experimental Setup
Our framework is built upon Wan2.1-I2V-14B-480P [wan2025wan], a publicly available image-to-video Diffusion Transformer with 40 layers. We adhere to a training-free paradigm, requiring no optimization or fine-tuning of the pre-trained diffusion model. For inversion, we utilize RF-Solver [wang2024taming] with 50 steps to map the source video into a noise latent, and the subsequent editing also uses 50 sampling steps. We set the timestep threshold $\tau$ to 0.5 (i.e., the mechanism is adopted for the first 50% of timesteps) and a guidance scale of 3.0.
@@
We evaluate our method on the Unic-Benchmark [ye2025unic]. For first-frame editing, we use AnyDoor [chen2024anydoor] for insertion, InsertAnything [song2025insert] for swapping, and MagicQuill [liu2025magicquill] for deletion. We compare against baselines including training-free AnyV2V [ku2024anyv2v], our adapted AnyV2V-DiT, and training-based VACE [chen2024vace] and I2VEdit [ouyang2024i2vedit]. For evaluation, we measure task-specific fidelity (CLIP-score, DINO-score), background preservation (PSNR, SSIM), and overall video quality using metrics from VBench [huang2024vbench] (e.g., consistency, smoothness, dynamic and aesthetic quality).
@@ 5 Conclusion
  • Additional Qualitative Comparisons. We present more visual comparisons against some state-of-the-art video editing techniques methods, specifically Unic [ye2025unic] and Pika [PikaLabs2025Pika2], to further highlight the superiority of ContextFlow.
@@ 6.1 Experimental Setup and Computational Cost
This section expands on the implementation details mentioned in the main paper. Our framework is implemented in Python using PyTorch. All experiments were conducted on a server with two NVIDIA A800 (80GB) GPUs. We build our training-free framework upon the publicly available Wan2.1-I2V-14B-480P [wan2025wan] model. The dual-path generation process, which is central to our Adaptive Context Enrichment mechanism, requires approximately 120GB of VRAM in total ( 60GB per GPU). For a typical 81-frame video at 480p resolution, the dual-path generation process takes approximately 25 minutes. As our method propagates an edit from the first frame, we utilize off-the-shelf image editing tools to generate the initial edited frame $I_{edit}$. For the experiments presented, we use AnyDoor [chen2024anydoor] for object insertion, InsertAnything [song2025insert] for object swapping, and MagicQuill [liu2025magicquill] for object deletion.
@@ 6.2 Further Details on Guidance Responsiveness Analysis
Expanding on the Guidance Responsiveness analysis presented in Figure 5 of the main paper, we offer details on its robustness. The $GR$ curves were derived by averaging results from a diverse set of 10 videos for each task, including in-domain videos from the Unic-Benchmark [ye2025unic] and out-of-domain videos from sources like the DAVIS dataset [ponttuset20182017davischallengevideo].
@@ 6.3 Evaluation Protocol Details
To measure how well the identity of the reference object is preserved, we compute frame-wise similarity. For CLIP-I, we use the pre-trained clip-vit-large-patch14 [Radford2021Learning] model from OpenAI. For DINO-I, we use the dinov2-giant [oquab2023dinov2] model from Facebook. The process is as follows: (1) we extract the image feature of the reference object; (2) for each frame in the generated video, we extract its image feature; (3) we compute the cosine similarity between the reference feature and each frame’s feature; (4) the final score is the average similarity across all frames.
@@
These metrics are computed exclusively on the unedited regions of the video to measure background stability. The unedited area is defined by a mask. For object swapping and deletion tasks on the Unic-Benchmark, we use the ground-truth masks provided with the dataset. For the object insertion task, where no prior object exists, we generate a mask for the newly inserted object using Grounding-SAM [ren2024grounded], and the background is defined as the area outside this mask. The final PSNR and SSIM scores are averaged over all frames.
@@
For the reader’s convenience, we briefly summarize the calculation methods for the VBench metrics used in our paper, drawing from the descriptions in the official VBench suite [huang2024vbench].
  • Overall Consistency. This metric evaluates the overall video-text consistency, reflecting both semantic and style alignment. It is computed using ViCLIP [Wang2023InternVid] as an aiding metric, where the model assesses the video against text prompts that contain different semantics and styles.
  • Motion Smoothness. To evaluate temporal quality beyond simple appearance consistency, this metric assesses whether the motion in the generated video is smooth and follows the physical laws of the real world. It utilizes the motion priors from a video frame interpolation model [Li2023AMT] to quantify the smoothness of the generated motions.
  • Dynamic Degree. Complementing smoothness, this metric measures the degree of dynamics (i.e., the presence of large motions) to ensure that completely static videos are not unfairly favored in temporal quality scores. It uses the RAFT [Teed2020RAFT] model to estimate the degree of dynamics in synthesized videos.
  • Aesthetic Quality. This metric evaluates the frame-wise artistic and beauty value perceived by humans. It employs the LAION [LAION-AI] aesthetic predictor to assess aspects such as the layout, the richness and harmony of colors, photo-realism, naturalness, and the overall artistic quality of the video frames.
@@ 7 Additional Qualitative Comparisons
To further situate ContextFlow’s performance, we provide qualitative comparisons against more strong, contemporary methods, for example Unic [ye2025unic] and commercial training-based tools like Pika  [PikaLabs2025Pika2] and Kling  [KlingAI_Global]. As shown in Figure 7 - Figure 14 , ContextFlow consistently demonstrates superior performance in both identity preservation and motion consistency.
@@ 8 Visual Analysis for Ablation Studies
Figure 20 and Figure 21 validates our data-driven layer selection strategy. Injecting information into all layers will lead to an excessive amount of information from the reconstruction path being introduced into the editing path, making it difficult to maintain the edited objects and causing the generated video to unconsciously lean towards the original video. Injecting guidance into the four least-responsive layers yields poor results, similar to providing weak guidance. Using a U-Net-based heuristic (from AnyV2V [ku2024anyv2v]) in the DiT model, which in a sense amounts to randomly selecting certain layers within DiT for injection, still leads to suboptimal results. Due to the inability of the injected layers to solve substantive problems, it often leads to stiff interaction between new objects and the background, and even unconscious emergence of unreasonable phenomena (such as the chimpanzee in Figure 21 strangely growing a tail). This visually confirms that our Guidance Responsiveness metric correctly identifies the most impactful layers for intervention.
@@
Figure 13: Visualization of Qualitative Comparison. Target prompt: “A hand-held white plastic watering can is used to water a vibrant collection of pink and purple flowers, including geraniums and possibly petunias, in a white planter.” The target to be deleted in this video is the railing in the lower left corner. Our method has high fidelity for the unedited areas, and its effect is better than that of Kling  [KlingAI_Global].
@@
Figure 20: Visualization of Ablation on Layer Selection. Target prompt: “A Pikachu is floating on the sea.” Injection on all layers and Injection on 4 least responsive layers will cause Pikachu to disappear quickly in subsequent frames or fail to display actions correctly according to the target prompt. Injection selection follow AnyV2V [ku2024anyv2v] will produce an extremely stiff and weird “floating” effect that is highly unnatural, along with an odd smile not required by the prompt. Please zoom in for a closer check.
Figure 21: Visualization of Ablation on Layer Selection. Target prompt: “A red poppy flower surrounded by purple flowers. A large gorilla is gently trying to touch the red poppy flower.” Injection on all layers and Injection on 4 least responsive layers will cause the inserted object to disappear or become translucent. Injection selection follow AnyV2V [ku2024anyv2v] will once again result in the strange form of the gorilla: an oddly appearing tail (within the red circle). Please zoom in for a closer check.